Grade 11Physical Chemistry

The density of 3.6M H2SO4 which is 29% by mass is g/ml is

Vidya
8 Years agoGrade 11

1 Answer

Rohit
8 Years ago
M= (10×X×d)/molar mass
Here X is concentration given in percentage.
Hence , d= (3.6×98)/(10×29)= 1.21.
Hope it helps
Regards 
ROHIT...